# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ 6.3.1 Proof of Lemma 12
Only the proof of Lemma 12 is left. We will make use of the following results in our proof.
Proposition 13 (Theorem 1 in [25]).
@@
Proposition 14 (Lemma D.8 in [18]; reformulated).
Let $G$ be a graph and $H$ be a subgraph of $G$. Let $C$ and $C^{\prime}$ be two 2EC components of $H$ such that $C$ is a $\mathcal{C}_{5}$ and $C^{\prime}$ is a $\mathcal{C}_{4}$, and there is a $3$-matching between $C$ and $C^{\prime}$ in $G$. Then, for any two vertices $x,y\in V(C)$ there exists a set $F$ of $9$ edges in $G$ such that $F\cup\{xy\}$ induces a 2EC spanning subgraph of $V(C)\cup V(C^{\prime})$ in $G$.
@@
Let $S$ be a non-trivial segment of ^$\bm{\hat{}}{G}_{H}$ and $A$ be a node in $S$. Then, there exists a matching of size 3 in $G$ between $V(C_{A})$ and $V(S)\setminus V(C_{A})$.
@@
If $A$ is not a cut node in ^$\bm{\hat{}}{G}_{H}$, then, by the definition of a segment, there are no edges between $V(C_{A})$ and the vertices of components which are not contained in $S$. Thus, in this case the claim follows from Lemma 2 using $(V(C_{A}),V(G)\setminus V(C_{A}))$ as the partition of $V(G)$. Otherwise, $A$ is a cut node in ^$\bm{\hat{}}{G}_{H}$, which implies $C_{A}$ is a separator of $G$ with the vertices of $V(S)\setminus V(C_{A})$ on one side of the separator (possibly along with some other vertices of $V(G)$). Consequently, in this case, our claim follows by applying Lemma 2 on a partition of $V(G)$ into two parts, where one part contains the vertices of $V(S)\setminus V(C_{A})$ (one side of the separator mentioned above) and the other part contains the vertices from $V(C_{A})$ such that the only edges crossing the two parts are between $V(S)\setminus V(C_{A})$ and $V(C_{A})$. ∎
We are now ready to prove Lemma 12.
Proof of Lemma 12.
@@
We are given a non-trivial segment $S$ of $G_{H}$ that contains a huge node $L$ and a node $A$ such that $C_{A}$ is either a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$. We need to construct in polynomial time a cycle $K$ in $G_{H}$ through $L$ and $A$, such that the two cycle edges incident on $C_{A}$ are incident on two vertices $u$ and $v$ in $C_{A}$ such that either (a) we can find a $u-v$ Hamiltonian path in $G[V(C_{A})]$ or (b) we can find a node $D$ outside the cycle $K$ and a set of edges $F$ disjoint to $E(K)$ such that $|F|=|E(C_{A})|+|E(C_{D})|$ and $F\cup\{uv\}$ forms a 2-ECSS of $G[V(C_{A})\cup V(C_{D})]$.
@@
In the proof we will show the existence of the cycle $K$ and a desired Hamiltonian path or a desired $D$ and $F$ with $|F|=9$. The polynomial-time construction for these objects then easily follow; $K$ from the algorithm in Proposition 13 by feeding it all possible $4$ edges, $2$ incident on $L$ and $2$ incident on $A$, and then the Hamiltonian path or $F$ can be found by brute-force enumeration over a constant domain.
@@
Let the cycle $C_{A}$ be labelled as $v_{1}-v_{2}-v_{3}-v_{4}-v_{1}$ if it is a $\mathcal{C}_{4}$ or $v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{1}$ if is is a $\mathcal{C}_{5}$. Since $S$ is a segment, it is 2-node-connected and, thus, there exist two internally node-disjoint paths $P_{1}$ and $P_{2}$ in $S$ between $L$ and $A$. Moreover, Lemma 15 implies that at least three vertices of $V(C_{A})$ have incident edges in $G$ from the vertices in $S$. We consider two exhaustive cases based on which set of vertices of $C_{A}$ has such incident edges: Either (Case 1) each of $v_{1},v_{2},v_{3}$ or (Case 2) each of $v_{1},v_{2},v_{4}$ has an incident edge from the vertices of $S$ that are not in $C_{A}$. Observe up to symmetry these are the only cases, and for $C_{A}$ being a $\mathcal{C}_{4}$ considering only Case 1 is sufficient. Thus, in Case 2, we will assume $C_{A}$ is a $\mathcal{C}_{5}$.
Case 1: We have two subcases.
@@
Case 1.1: There exists paths $P_{1}$ and $P_{2}$ that are incident to vertices $u$ and $v$ of $C_{A}$ such that there exists a Hamiltonian Path between $u$ and $v$ in $G[V(C_{A})]$. Then, $K=P_{1}\cup P_{2}$ forms a cycle in ^$\bm{\hat{}}{G}_{H}$ through $A$ and $L$ as $P_{1}$ and $P_{2}$ are internally node-disjoint, which satisfies the requirements of (a).
@@
Case 1.2: Otherwise, if we are not in Case 1.1, there must exist paths $P_{1}$ and $P_{2}$ such that $P_{1},P_{2}$ are incident to $V(C_{A})$ at vertices $u_{1},u_{2}\in V(C_{A})$, respectively ($u_{1}=u_{2}$ is possible), and there exist $w_{i}\in\{v_{1},v_{2},v_{3}\}$ adjacent to $u_{i}$ in $C_{A}$ for $i\in\{1,2\}$. This observation uses that $C_{A}$ is a cycle of at most 5 vertices. Consider two subcases:

### v2

@@ 6.3.1 Proof of Lemma 11
Only the proof of Lemma˜11 is left. We will make use of the following results in our proof.
Proposition 12 (Theorem 1 in [Kaw08]).
@@
Proposition 13 (Lemma D.8 in [GGA23]; reformulated).
Let $G$ be a graph and $H$ be a subgraph of $G$. Let $C$ and $C^{\prime}$ be two 2EC components of $H$ such that $C$ is a $\mathcal{C}_{5}$ and $C^{\prime}$ is a $\mathcal{C}_{4}$, and there is a $3$-matching between $C$ and $C^{\prime}$ in $G$. Then, for any two distinct vertices $x,y\in V(C)$ there exists a set $F$ of $9$ edges in $G$ such that $F\cup\{xy\}$ induces a 2-ECSS on $V(C)\cup V(C^{\prime})$ in $G$.
@@
Let $S$ be a non-trivial segment of $\bm{\hat{}}{G}_{H}$ and $A$ be a node in $S$. Then, there exists a matching of size 3 in $G$ between $V(C_{A})$ and $V(S)\setminus V(C_{A})$.
@@
If $A$ is not a cut node in $\bm{\hat{}}{G}_{H}$, then, by the definition of a segment, there are no edges between $V(C_{A})$ and the vertices of components which are not contained in $S$. Thus, in this case the claim follows from Lemma˜2 using $(V(C_{A}),V(G)\setminus V(C_{A}))$ as the partition of $V(G)$. Otherwise, $A$ is a cut node in $\bm{\hat{}}{G}_{H}$, which implies $C_{A}$ is a separator of $G$ with the vertices of $V(S)\setminus V(C_{A})$ contained on the same side of the separator. Consequently, in this case, our claim follows by applying Lemma˜2 on a partition of $V(G)$ into two parts: one side of the separator that contains the vertices of $V(S)\setminus V(C_{A})$ and the other part contains the vertices from $V(C_{A})$ along with the other side of the separator such that the only edges crossing the two parts are between $V(S)\setminus V(C_{A})$ and $V(C_{A})$. ∎
We are now ready to prove Lemma˜11.
Proof of Lemma˜11.
@@
We are given a non-trivial segment $S$ of $G_{H}$ that contains a huge node $L$ and a node $A$ such that $C_{A}$ is either a $\mathcal{C}_{4}$ or a $\mathcal{C}_{5}$. We need to construct in polynomial time a cycle $K$ in $G_{H}$ through $L$ and $A$, such that the two cycle edges incident on $C_{A}$ are incident on two distinct vertices $u$ and $v$ in $C_{A}$ such that either (a) we can find a uvu-v Hamiltonian path in $G[V(C_{A})]$ or (b) we can find a node $D$ outside the cycle $K$ and a set of edges $F$ disjoint to $E(K)$ such that $|F|=|E(C_{A})|+|E(C_{D})|$ and $F\cup\{uv\}$ forms a 2-ECSS of $G[V(C_{A})\cup V(C_{D})]$.
@@
In the proof, we will show the existence of the cycle $K$ and a desired Hamiltonian path or a desired $D$ and $F$ with $|F|=9$. The polynomial-time constructions for these objects then easily follow; $K$ from the algorithm in Proposition˜12 by feeding it all possible $4$ edges, $2$ incident on $L$ and $2$ incident on $A$, and then the Hamiltonian path or $F$ can be found by brute-force enumeration over a constant domain.
@@
The cycle $C_{A}$ will be labeled as $v_{1}-v_{2}-v_{3}-v_{4}-v_{1}$ if it is a $\mathcal{C}_{4}$ or $v_{1}-v_{2}-v_{3}-v_{4}-v_{5}-v_{1}$ if it is a $\mathcal{C}_{5}$. Since $S$ is a segment containing $L$ and $A$, it is 2-node-connected, and thus, there exist two internally node-disjoint paths in $S$ between $L$ and $A$. Moreover, Lemma˜14 implies that at least three distinct vertices of $V(C_{A})$ have incident edges in $G$ from three distinct vertices in $V(S)\setminus V(C_{A})$. We consider a general condition first that allows us to find certain cycles in $S$ between the nodes $L$ and $A$. This condition will show up multiple times in our analysis.
Condition ($\star$) Let $P_{1}$ and $P_{2}$ be two internally node-disjoint paths from $L$ to $A$ that are incident on vertices $x$ and $y$ of $C_{A}$, respectively. Additionally, let $z$ be a vertex in $C_{A}$ that is adjacent to some node $A^{\prime}\neq A$ in the segment $S$. We can now find the following cycles in $S$ that pass through the nodes $L$ and $A$. Note that $x$, $y$, and $z$ may not be all distinct.
  • Since $A^{\prime}$ is part of the segment $S$, there are two internally node-disjoint paths from $A^{\prime}$ to $L$. At least one of these paths will not include the node $A$, say $P^{\prime}$.
  • $(\star 1$) If $P^{\prime}$ is node-disjoint from $P_{1}$, except at $L$, then there is a cycle that intersects $C_{A}$ at $x$ and $z$: $z\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $$}}$}}A^{\prime}\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}$}}$}}L\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P_{1}$}}$}}x$. Similarly, if $P^{\prime}$ is node-disjoint from $P_{2}$, except at $L$, then there is a cycle that intersects $C_{A}$ at $y$ and $z$.
  • Otherwise, if $P^{\prime}$ is not node-disjoint from $P_{1}$, let $B$ be the first node when traversing $P^{\prime}$ from $A^{\prime}$ to $L$ that intersects $P_{1}$ or $P_{2}$. Note that $B$ cannot be on both $P_{1}$ and $P_{2}$ as they are internally node-disjoint.
  • $(\star 2)$ If $B$ is on $P_{1}$, then we have the cycle that intersects $C_{A}$ at $y$ and $z$: $z\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $$}}$}}A^{\prime}\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}$}}$}}B\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P_{1}$}}$}}L\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P_{2}$}}$}}y$. If $B$ is on $P_{2}$, similarly we get a cycle that intersects $C_{A}$ at $x$ and $z$.
  • No matter which case we are in, there is always a cycle intersecting $C_{A}$ at $z$ and ($x$ or $y$).
As mentioned above, Lemma˜14 implies that at least three distinct vertices of $V(C_{A})$ have incident edges in $G$ from three distinct vertices in $V(S)\setminus V(C_{A})$. We consider two exhaustive cases based on which set of vertices of $C_{A}$ has such incident edges: Either (Case 1) each of $v_{1},v_{2},v_{3}$ or (Case 2) each of $v_{1},v_{3},v_{4}$ has an incident edge from the vertices of $V(S)\setminus V(C_{A})$. Observe up to symmetry these are the only cases, and for $C_{A}$ being a $\mathcal{C}_{4}$ considering only Case 1 is sufficient. Thus, in Case 2, we will assume $C_{A}$ is a $\mathcal{C}_{5}$.
Case 1: $v_{1}$, $v_{2}$, and $v_{3}$ have incident edges from vertices of $V(S)\setminus V(C_{A})$.
@@
Case 1.1: There exists two internally node-disjoint paths $P_{1}$ and $P_{2}$ from $A$ to $L$ that are incident to distinct vertices $u$ and $v$ of $C_{A}$ such that there exists a Hamiltonian Path between $u$ and $v$ in $G[V(C_{A})]$. Then, $K=P_{1}\cup P_{2}$ forms a cycle in $\bm{\hat{}}{G}_{H}$ through $A$ and $L$ as $P_{1}$ and $P_{2}$ are internally node-disjoint, which satisfies the requirements of (a) in the lemma statement.
@@
Case 1.2: Otherwise, if we are not in Case 1.1, there must exist two internally node-disjoint paths $P_{1}$ and $P_{2}$ from $A$ to $L$ such that $P_{1},P_{2}$ are incident to $V(C_{A})$ at vertices $u_{1},u_{2}\in V(C_{A})$, respectively ($u_{1}=u_{2}$ is possible, but $u_{1}$ is not a neighbor of $u_{2}$). Since, $C_{A}$ is a cycle consisting of at most 5 vertices, we can always find $w_{i}\in\{v_{1},v_{2},v_{3}\}$ adjacent to $u_{i}$ in $C_{A}$ for each $i\in\{1,2\}$. If possible, set $w_{1}=w_{2}$. In particular, if $u_{1}=u_{2}$, we can always set $w_{1}=w_{2}\in\{v_{1},v_{2},v_{3}\}$. Furthermore, if $C_{A}=\mathcal{C}_{4}$, $u_{1}\neq u_{2}$, and $u_{1}$ not adjacent to $u_{2}$ (otherwise we would be in Case 1.1) implies they are diagonally opposite, and we can again set $w_{1}=w_{2}\in\{v_{1},v_{2},v_{3}\}$. Consider two subcases:
Case 1.2.1: $u_{1}=u_{2}$ or $w_{1}=w_{2}$. This implies that both $u_{1}$ and $u_{2}$ are adjacent to $w_{1}$. Since $w_{1}\in\{v_{1},v_{2},v_{3}\}$ it is adjacent to a node $A^{\prime}\neq A$ in $S$. We now apply Condition ($\star$) by setting $x=u_{1}$, $y=v_{1}$, and $z=w_{1}$ to find a cycle through $L$ and $A$ that intersects $C_{A}$ at $w_{1}$ and ($u_{1}$ or $u_{2}$), and since $w_{1}$ is adjacent to both $u_{1}$ and $u_{2}$, we can find a desired Hamiltonian path in $G[V(C_{A})]$ satisfying condition (a) of the lemma statement.
Case 1.2.2: $u_{1}\neq u_{2}$ and $w_{1}\neq w_{2}$. Since we are not in the previous case, we only need to consider $C_{A}=C_{5}$ and $\mathsf{dist_{C_{A}}}(u_{1}=u_{2})=2$. Observe that in this case we can always set $w_{1}$ and $w_{2}$ to be adjacent vertices. Recall that $w_{i}\in\{v_{1},v_{2},v_{3}\}$ for each $i\in\{1,2\}$ and since we are in Case 1, both $w_{1}$ and $w_{2}$ have incident edges from vertices of $V(S)\setminus V(C_{A})$. Let $A_{i}\neq A$ be the node in $S$ that is adjacent to $w_{i}$, for each $i\in\{1,2\}$. Since $S$ is 2-node connected, there must be a path $P^{\prime}_{i}$ in $S$ from $A_{i}$ to $L$ which does not contain $A$, for $i\in\{1,2\}$.
If $P^{\prime}_{1}$ is node-disjoint from $P_{1}$ or if $P^{\prime}_{2}$ is node-disjoint from $P_{2}$, then we can find a desired cycle using Condition ($\star 1$) that intersects $C_{A}$ at adjacent vertices $u_{i}$ and $w_{i}$ for some $i\in\{1,2\}$ which have a Hamiltonian path between them in $C_{A}$. Otherwise, for each $i\in\{1,2\}$, let $B_{i}$ be the first node on $P^{\prime}_{i}$ when traversing $P^{\prime}_{i}$ from $A_{i}$ to $L$ that is also on either $P_{1}$ or on $P_{2}$. Now if $B_{1}$ is on $P_{2}$ or $B_{2}$ is on $P_{1}$, we reach the same conclusion using Condition ($\star 2$).
Otherwise, the paths $P^{\prime}_{1}$ and $P^{\prime}_{2}$ are such that $B_{1}$ is on $P_{1}$ and $B_{2}$ is on $P_{2}$, and, thus, the subpaths $A_{1}\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}_{1}$}}$}}B_{1}$ and $A_{2}\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}_{2}$}}$}}B_{2}$ must be node-disjoint; if not, let $t$ be the first node on $A_{1}\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}_{1}$}}$}}B_{1}$ that is also on $A_{2}\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}_{2}$}}$}}B_{2}$. We will modify $P^{\prime}_{1}$ to be $A_{1}\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}_{1}$}}$}}t\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}_{2}$}}$}}B_{2}\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}_{2}$}}$}}L$, which now has its first intersecting node with $P_{1}$ or $P_{2}$ to be $B_{2}$, which is on $P_{2}$, and we are back to the case considered above. But then there exist two-node disjoint paths from $A$ to $L$ in $\bm{\hat{}}{G}_{H}$ that are incident to $w_{1}$ and $w_{2}$: $L\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P_{i}$}}$}}B_{i}\mathrel{\raisebox{-1.3pt}{$\xrightarrow{\mbox{ $P^{\prime}_{i}$}}$}}A_{i}-w_{i}$ for $i\in\{1,2\}$, and there is a Hamiltonian Path from $w_{1}$ to $w_{2}$ in $G[V(C_{A})]$ as $w_{1}$ and $w_{2}$ are adjacent, meaning that we are actually in Case 1.1.
Case 2: $v_{1}$, $v_{3}$, and $v_{4}$ have incident edges from vertices of $V(S)\setminus V(C_{A})$. As discussed above, this case can only apply if $A$ corresponds to a $\mathcal{C}_{5}$. Further, note that neither $v_{2}$ nor $v_{5}$ can have an incident edge to a node in $S$ other than $A$, as otherwise we are also in Case 1 by simply relabeling the $v_{i}$’s cyclically so that $v_{1},v_{2},v_{3}$ have such incident edges. We have two subcases based on whether $v_{2}$ or $v_{5}$ has an incident edge to a component outside of segment $S$.
Case 2.1: Neither $v_{2}$ nor $v_{5}$ has incident edges to nodes of $\bm{\hat{}}{G}_{H}$ other than $A$. Then, there must be an edge between $v_{2}$ and $v_{5}$ in $G$, as otherwise $C_{A}$ is contractible: This is true since then the set of edges incident to $v_{2}$ and the set of edges incident to $v_{5}$ are all inside $G[V(C_{A})]$ and mutually disjoint. Since any 2-ECSS of $G$ must include 2 edges incident on each vertex, it has to contain at least 4 edges from $G[V(C_{A})]$, and hence $C_{A}$ is contractible. Like in Case 1.1, if there exists internally node-disjoint paths $P_{1}$ and $P_{2}$ that are incident to vertices $u$ and $v$ in $C_{A}$ such that there exists a Hamiltonian Path between $u$ and $v$ in $G[V(C_{A})]$, then $K=P_{1}\cup P_{2}$ forms the desired cycle in $\bm{\hat{}}{G}_{H}$ through $A$ and $L$ satisfying the requirements of Condition (a) in the lemma statement. As a result, observe that both $P_{1}$ and $P_{2}$ must be incident to the same vertex $v\in\{v_{1},v_{3},v_{4}\}$. This is true since neither $P_{1}$ nor $P_{2}$ can be incident on $v_{2}$ or $v_{5}$. Furthermore, the edge $v_{2}v_{5}$ exists helping in forming the required Hamiltonian paths; a straightforward check. For example, $v_{1}-v_{2}-v_{5}-v_{4}-v_{3}$ is a Hamiltonian $v_{1}$-$v_{3}$-path. Now, If $v=v_{1}$ (a similar argument works if $v=v_{3}$ or $v=v_{4}$), let $A^{\prime}\neq A$ be the node adjacent to $v_{3}$ ($v_{1}$ if $v=v_{3}$ or $v=v_{4}$). Now from Condition ($\star$), by setting $x=y=v$, and $z=v_{3}$, we get the desired cycle passing through $A$ and $L$ that intersects $C_{A}$ at $v_{1}$ and $v_{3}$ having a Hamiltonian path between them, satisfying Condition (a) of the lemma statement.
Case 2.2: There is an edge between $v_{2}$ (or $v_{5}$ by symmetry) and a node $D$ of $\bm{\hat{}}{G}_{H}$ that is a trivial segment $S^{\prime}$ or lies in a non-trivial segment $S^{\prime}$, and $S^{\prime}\neq S$. Note that $A$ must be a cut node in $\bm{\hat{}}{G}_{H}$ in this case. Observe that $|V(S)|-|V(C_{A})|\geq 10$, because $S$ contains the node $L$ ($C_{L}$ is huge with 10 vertices) other than $A$. If $|V(S^{\prime})\cup V(C_{A})|\geq 10$, Lemma˜3 guarantees the existence of a $4$-matching $M_{4}$ between $V(S)\setminus V(C_{A})$ and $V(C_{A})$ (as $C_{A}$ is a separator, $V(G)$ can be partitioned in 2 parts with $V(C_{A})\cup V(S^{\prime})$ on one side and $V(S)\setminus V(C_{A})$ on the other, so that the only edges crossing are between $V(C_{A})$ and $V(S)\setminus V(C_{A})$). But this means that we are actually in Case 1 as some three consecutive vertices of $C_{A}$ must be incident to the other nodes of $S$.
Thus, the only possibility left is $|V(S^{\prime})\cup V(C_{A})|\leq 9$. Since $A$ corresponds to a $\mathcal{C}_{5}$ as we are in Case 2, we conclude that $S^{\prime}$ is a trivial segment that corresponds to a single node $D$ where $C_{D}$ is a $\mathcal{C}_{4}$. Note that $|E(C_{A})|+|E(C_{D})|=9$. Since $A$ is a cut node and $D$ is a trivial segment, Proposition˜7 implies that there exists a $3$-matching $M_{3}$ between $V(C_{A})$ and $V(C_{D})$ in $G$. Therefore, Proposition˜13 is applicable to $A$ and $D$ with any two distinct vertices $u$ and $v$ in $C_{A}$, which implies that there exists a set $F$ of edges of $G$ with $|F|=9$ such that $F\cup\{uv\}$ induces a 2-ECSS of $G[V(C_{A})\cup V(C_{D})]$. If $P_{1}$ and $P_{2}$, the two internally node-disjoint paths between $L$ and $A$, are incident on two distinct vertices in $C_{A}$, then we set those two distinct vertices to $u$ and $v$, set $K:=P_{1}\cup P_{2}$, which together with $F$ satisfies the Condition (b) in the lemma statement. Otherwise, $P_{1}$ and $P_{2}$ are incident to the same vertex of $V(C_{A})$, which we call $u$, and let $A^{\prime}\neq A$ be a node adjacent to $v$, where $v\in\{v_{1},v_{3},v_{4}\}\setminus\{u\}$. Now from Condition ($\star$), there exists a cycle in $S$ passing through $A$ and $L$ that intersects $C_{A}$ at $u$ and $v$. This cycle together with $F$ satisfies Condition (b) in the lemma statement. ∎
7 Conclusion
In this work, we presented a new approach to the 2-edge-connected spanning subgraph (2-ECSS) problem, achieving an improved polynomial-time approximation ratio of $\nicefrac{{5}}{{4}}+\varepsilon$. This result follows a long line of research efforts aimed at improving the approximation ratio, building on prior methods while simplifying the complexity of the algorithm. Our method significantly streamlined the gluing step, which has traditionally been one of the most complex parts of such algorithms. By introducing a preprocessing phase to eliminate large 3-vertex cuts and leveraging a novel 4-matching lemma, we modularized the solution into manageable components. This led to a method that is not only simpler but also more intuitive than previous approaches.
The modularity and clarity of our approach opens new avenues for future research, providing a solid foundation for adaptations and further advancements. The combination of the 4-matching lemma and the Pac-Man-inspired gluing strategy we proposed could also prove useful in tackling other network design problems that are hindered by complex structural challenges.
Appendix A Reduction to Structured Graphs
This section is dedicated to the proof of Definition˜5, which we first restate.